A 5.00L flask contains 0.600g of O2 at a temperature of 22oC. What is the pressure in atm inside the flask?
1
Expert's answer
2020-10-16T08:55:25-0400

"n=\frac{0.6}{32}=0.0188"


"T=273+22=295K"

Applying PV=nRT

"P\times5=0.0188\times0.082\times295"

"P=0.09 \ atm"